OpenStudy (anonymous):

graph the line that has a y intercept (0,-1) and a slope of -1/3. Label two points

OpenStudy (wwb00):

Let's work together. Ok?

OpenStudy (wwb00):

The Y intercept is the point where the line crosses the y axis. Thus, this must be a point on your line.

OpenStudy (wwb00):

To find another point follow the slope downwards, your slope is -1/3, and according to rise over run, every three units to the right, your line moves down one unit. Thus you should be able to graph another point of the line.

OpenStudy (wwb00):

Do you understand how to find another point of the line?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-y=m(x-x)

OpenStudy (texaschic101):

a y intercept of (0,-1) and a slope of -1/3. y = -1/3x - 1 .... this is your equation of that line. now if you want to find more points, just start subbing in numbers for x to find y y = -1/3x - 1....when x = 0 y = -1/3(0) - 1 y = -1.....so (0,-1) is on the line and it is also the y intercept. you can either sub in numbers for x and finding y...or...you could just find the x intercept by subbing in 0 for y. y = -1/3x - 1 0 = -1/3x - 1 1/3x = -1 x = -1/(1/3) x = -1 * 3 x = -3....so another points is (-3,0)...this is also your x intercept